### v3

@@
To find families for the more than 100,000 children in need of adoptive placements, most United States child welfare agencies have employed a family-driven search strategy in which prospective families respond to announcements made by the agency. However, some agencies have switched to a caseworker-driven search strategy in which the caseworker directly contacts families recommended for a child. We introduce a novel search-and-matching model to capture essential aspects of the adoption process and compare the two approaches through a game-theoretical analysis. We show that the search equilibria induced by threshold strategies form a lattice structure under either approach. Our main theoretical result establishes that the equilibrium outcomes in family-driven search can never Pareto dominate the outcomes in caseworker-driven search, but it is possible that each caseworker-driven search outcome Pareto dominates all family-driven search outcomes. We also find that when families are sufficiently impatient, caseworker-driven search is better for all children. We illustrate numerically that most agents are better off under caseworker-driven search for a wide range of parameters. Finally, we provide empirical evidence from an agency that switched to caseworker-driven search and achieved a three-year adoption probability that outperformed a statewide benchmark by 17%.
@@ 1 Introduction
Child welfare systems worldwide face the challenge of finding families for children in need of adoption. For example, the United States foster care system serves over 600,000 children annually, with approximately 400,000 children in temporary foster care arrangements at any time. While the goal for most of these children is to reunite them with their parents or close relatives, roughly 120,000 children are waiting for permanent adoptive placements (Children’s Bureau 2022). Finding an adoptive family for these children has become a public policy priority due to high levels of incarceration, homelessness, unemployment, and teen pregnancy observed in the population of children “aging out” of the child welfare system without a permanent family relationship (Barth 1990, Triseliotis 2002, Kushel et al. 2007). In this paper, we study the search and matching process for children waiting in the child welfare system for adoptive placement and compare two prominent search methods that differ in whether families or the children’s caseworkers drive the search process. We provide structural insights by analytically and numerically analyzing a game-theoretic model and empirically validate our findings by studying outcomes achieved by a Florida child welfare agency that switched its search strategy in 2018.
@@
The search for an adoptive placement officially begins once a judge issues a termination of parental rights order. A caseworker represents the child’s interests throughout this process to find an adoptive family if adoption by a relative or foster care family is not possible. Identifying a family willing to adopt the child and capable of caring for the child can be a difficult task, and the challenge varies greatly according to the child’s demographic characteristics and special needs. While relatively little research has studied best practices for search operations — either through empirical studies of how caseworkers find families or prescriptive studies for how search should be conducted — states invest significant resources in trying to help the most vulnerable children find permanency; for example, in Florida’s 2022-2023 fiscal year, the state’s child welfare agencies spent \$28 million just to recruit adoptive homes and improve adoption outcomes for a population of children eligible for adoption that numbered 3,787 on July 1, 2022 (Florida Department of Children and Families 2023).
@@
Different states, counties, and agencies adhere to different paradigms for the practice of search and matching. Families first register with an adoption agency and provide a home study evaluation. In the predominant approach, caseworkers then announce children via email to a set of registered families, each of whom has the opportunity to express interest in a child. The caseworker receives these inquiries and works to identify the family that best fits the child’s needs. We label this approach family-driven search, as families direct the search process by expressing interest in available children. Hanna and McRoy (2011) shows how the child welfare literature focuses almost exclusively on this family-driven search process. One important downside of this search process is that some children may attract hundreds of interested families, all of which the caseworker has to consider simultaneously.
@@
These interactions can be time-consuming and emotionally stressful for all parties involved, and some states mandate that agencies engage with every family that has expressed interest. For example, Florida Administrative Code Rule 65C-16.003 states:
@@
We have also heard families share their frustration with the emotional and time costs that they incurred while participating in a family-driven search system. In a video shared on social media, one adoptive father described his journey that began as a prospective parent engaged with an adoption agency that announced available children and had families respond to express interest in particular children: [^1] https://www.linkedin.com/posts/adoption-share_by-inverting-the-outreach-for-adoption-matching-activity-7131265613947138048-YCM1
@@
In reaction to these challenges, some child welfare agencies have recently sought to improve outcomes by switching to caseworker-driven search (Riley 2019). In this approach, caseworkers sequentially contact specific prospective families to share details about the child. Caseworkers use their expertise to decide which families to contact based on the child’s and the families’ characteristics. Optionally, technological tools may also aid their decisions. This approach removes the burden of having to engage with a large number of families simultaneously and allows caseworkers to target compatible families for children with very specific needs.
@@
However, even though these are clear advantages of caseworker-driven search, it is important to note that changing the search paradigm may change the incentives of families to be interested in different types of children. For a match to occur in either discipline, both the caseworker and the prospective family need to agree. This can for example mean that if caseworker-driven search leads to a reduction in search costs for families, they may stop being interested in certain children, leaving those children unmatched. As such, it is unclear whether case-worker-driven search leads to more desirable outcomes for the entire population of children requiring adoptive placements.
@@
To assess how different search strategies affect outcomes, we pursue two complementary approaches in this paper. First, we analyze the strategic behavior of agents under both search disciplines in a game-theoretic search-and-matching model. Second, we provide empirical evidence for the performance of caseworker-driven search compared to classic search approaches by studying outcomes of a real-world child welfare agency.
@@
Our analytical contributions start with introducing a new search-and-matching model (Section 3), which captures critical features: First, as both approaches to adoption matching are inherently dynamic, we assume children and families (simply agents from now on) may enter and depart the system at any time. However, to keep the model tractable, we assume the distribution of agent types in the system remains stable over time. Second, we allow for uncertainty regarding whether a child-family pair would actually be compatible. Third, our model captures the heterogeneous preferences of agents, which is a key distinction between our paper and most earlier literature. While we focus specifically on analyzing adoption systems, our work is also relevant to more general search-and-matching theory. To the best of our knowledge, we are the first to formally analyze economic effects in the child welfare domain while taking heterogeneity of preferences into account.
@@
We perform a game-theoretic comparison of the two approaches within our model. In Section 4.2, we establish that equilibria are guaranteed to exist under both search technologies (i.e., family-driven search and caseworker-driven search). Furthermore, we find that equilibria form a lattice reminiscent of the structure of the set of stable matchings in standard two-sided matching markets. We then present our main theoretical result: Family-driven search equilibrium outcomes can never be Pareto improvements over caseworker-driven search equilibrium outcomes, but there are instances where each caseworker-driven search equilibrium outcome is a Pareto improvement over all family-driven search equilibrium outcomes (Section 5.1). This holds because caseworker-driven search can reduce wasted search efforts. Thus, agents can worry less about accumulating search costs when they express interest in a child. However, because of multiplicity of equilibria and the lattice structure over these equilibria, all children can be strictly better off in either system. The same holds for families. Even when both family-driven search and caseworker-driven search only admit a unique equilibrium each, an agent can be better off in either setting (Section 5.2). This may be surprising, given that caseworker-driven search reduces wasteful search efforts on both sides of the market. We therefore explore the conditions under which caseworker-driven search usually leads to more agents being better off. In Section 6.1, we show that all children will be better off in caseworker-driven search if families are sufficiently impatient. Furthermore, increasing family supply in family-driven search can have a negative effect on children’s utilities, but not in caseworker-driven search (Section 6.2). We find numerically that caseworker-driven search leads to more desirable outcomes for a wide variety of model parameter choices (Section 7).
@@
To supplement this analytical and numerical insight, we then analyze case-level data from a technology platform that a multi-county child welfare agency in Florida used for the majority of its search efforts since 2018 (Section 8). We compare the outcomes for hundreds of children in need of adoptive placements to a benchmark from statewide case-level data that accounts for a child’s demographic information and disabilities. Besides dramatically reducing the search efforts and skill required by caseworkers, we show that the probability of adoption within three years was 17% higher than the benchmark from statewide data.
@@ 2 Related Literature
Recently, there has been increased interest from market designers and operations researchers in the challenges faced by historically underserved and disadvantaged communities. Researchers have, for example, studied refugee resettlement (Andersson et al. 2018, Bansak et al. 2018, Delacrétaz et al. 2020), the improvement of teacher quality at disadvantaged schools (Combe et al. 2022), and the management of volunteer workforces for non-profit organizations (Berenguer et al. 2023).
@@
As part of this new focus on service systems for vulnerable populations, our paper contributes to the literature that studies child welfare systems from operations or market design perspectives — a challenge first articulated by (Spindler 1970). Slaugh et al. (2016) investigated how the Pennsylvania Statewide Adoption & Permanency Network could utilize a match recommendation tool to improve their process of matching children to prospective parents. Their spreadsheet-based tool can be seen as a simplified version of the previously mentioned data-driven software system. Robinson-Cortés (2019) worked with a foster care data set to analyze placements of children in foster homes. His model predicts that allowing placements across administrative regions would be beneficial for children. MacDonald (2019) studied a dynamic matching problem where children and families can either form reversible matches (foster placements) or irreversible matches (adoptions). In her model, children are heterogeneous in the sense that there are children with disabilities and children without disabilities, while families are homogeneous. Baccara et al. (2014) estimate families’ preferences over children available for adoption from a data set documenting the operations of adoption agencies. To the best of our knowledge, we are the first to formally analyze the economic effects of search and matching in the child welfare domain while taking heterogeneity of preferences into account. While we focus specifically on analyzing adoption systems, our work is also relevant for search-and-matching theory in general.
@@
Within the child welfare literature, relatively little research has investigated the effectiveness of search strategies for children in need of adoptive placements. Some research has reported positive impacts from intensive multi-faceted search efforts by case workers: Vandivere et al. (2015) show that children served by recruiters from the Wendy’s Wonderful Kids organization were 1.7 times as likely to have an adoptive placement than a control group in an experiment with over 1,000 children. In a similar context focusing on hard-to-place youth in New York, Feldman et al. (2016) showed that a program of enhanced casework improved outcomes for children. The program utilized a variety of channels to promote 88 children and search for them. The search methods in both experiments require extensive work from skilled caseworkers funded by grants, while the platform we study provides an example of technology assisting caseworkers. Avery et al. (2009) study national photolisting service AdoptUSKids and use a hazards model to show better outcomes for children based on activity on AdoptUsKids. However, photolistings have drawn increased scrutiny since the early 2000s; Roby and White (2010) describes risks for exploitation and bullying for children publicly listed online.
@@
Even though child adoption matching markets bear similarities to other two-sided matching markets such as centralized labor markets (Roth 1984, 1991) or ride-sharing platforms (Ma et al. 2020), there are important differences that necessitate new models and analyses. Adoption matching is inherently dynamic, and there is no centralized clearinghouse that determines final matches. Matches are only ever proposed, and both sides of the market have to invest search efforts to identify a match candidate. Purely random decentralized matching models have been widely studied under search frictions and homogeneous preferences (Eeckhout 1999, Shimer and Smith 2000, 2001, Atakan 2006) and more rarely under heterogeneous preferences (Adachi 2003). More recent studies have combined directional search rather than random search as an important feature (Lauermann and Nöldeke 2014, Lauermann et al. 2020, Cheremukhin et al. 2020).
@@
One matching market similar to adoption from foster care is online dating (Hitsch et al. 2010a, b, Lee and Niederle 2015, Rios et al. 2021, Kanoria and Saban 2021). However, online dating markets are completely decentralized despite their dynamic and recommendation-based features. Individuals in search of a romantic partner can, at any time, decide to browse a dating platform and reach out to other individuals who appeal to them. In contrast, the approaches we analyze in this paper follow a centralized protocol despite the dynamic decentralized search component: Caseworkers perform the search for a family on behalf of children, and they do so in approximately regular time intervals. Caseworkers therefore play an essential role throughout the process since they act as an intermediary to protect vulnerable children. Crucially, this introduces an asymmetry that we have not observed in any other previously studied matching market. As a consequence, we develop a new model that allows us to capture the features of the two different search technologies in one model.

### v4

@@
To find families for the more than 100,000 children in need of adoptive placements, most United States child welfare agencies have employed a family-driven search approach in which prospective families respond to announcements made by the agency. However, some agencies have switched to a caseworker-driven search approach in which the caseworker directly contacts families recommended for a child. We introduce a novel search-and-matching model that captures the key features of the adoption process and compare family-driven with caseworker-driven search in a game-theoretical framework. Under either approach, the equilibria are generated by threshold strategies and form a lattice structure. Our main theoretical finding then shows that no family-driven equilibrium can Pareto dominate any caseworker-driven outcome, whereas it is possible that each caseworker-driven equilibrium Pareto dominates every equilibrium attainable under family-driven search. We also find that when families are sufficiently impatient, caseworker-driven search is better for all children. We illustrate numerically that most agents are better off under caseworker-driven search for a wide range of parameters. Finally, we provide empirical evidence from an agency that switched to caseworker-driven search and achieved a three-year adoption probability that outperformed a statewide benchmark by 24%, as well as a statistically significant 27% improvement in adoption hazard rates.
@@ 1 Introduction
Child welfare systems worldwide face the challenge of finding families for children in need of adoption. For example, the United States foster care system serves over 525,000 children annually. While the goal for most of these children is to reunite them with their parents or relatives, 77,809 out of 343,077 children in foster care on September 30, 2023, were waiting for permanent adoptive placements (U.S. Administration for Children and Families 2024). Finding an adoptive family for these children has become a public policy priority due to high levels of incarceration, homelessness, unemployment, and teen pregnancy observed in the population of children “aging out” of the child welfare system without a permanent family relationship (Triseliotis 2002, Kushel et al. 2007, Gypen et al. 2017). In this paper, we study the search and matching process for children waiting in the child welfare system for adoptive placement and compare two prominent search methods that differ in whether families or the children’s caseworkers drive the search process. We provide structural insights by analytically and numerically analyzing a game-theoretic model, and empirically validate our findings by examining the outcomes achieved by a Florida child welfare agency that switched its search strategy in 2018.
@@
The search for an adoptive placement officially begins once a judge issues a termination of parental rights order. A caseworker represents the child’s interests throughout this process to find an adoptive family if adoption by a relative or foster care family is not possible. Identifying a family willing to adopt the child and capable of caring for the child can be a difficult task, and the challenge varies greatly according to the child’s demographic characteristics and special needs. While relatively little research has studied best practices for search operations — either through empirical studies of how caseworkers find families or prescriptive studies for how search should be conducted — states invest significant resources in trying to help the most vulnerable children find permanency; for example, the Florida Department of Children and Families (2022, 2024) reports spending over \$20 million annually to promote and support searches for approximately 3,800 children with a goal of adoption at any point in time.
@@
Different states, counties, and agencies adhere to different paradigms for the practice of search and matching. Families first register with an adoption agency and provide a home study evaluation. In the predominant approach, caseworkers then announce children via email to a set of registered families, each of whom has the opportunity to express interest in a child. The caseworker receives these inquiries and works to identify the family that best fits the child’s needs. We label this approach family-driven search, as families direct the search process by expressing interest in available children. Hanna and McRoy (2011) demonstrates how the child welfare literature has almost exclusively focused on what we view as family-driven search processes. One important downside of this search process is that some children may attract hundreds of interested families, all of which the caseworker has to consider simultaneously.
@@
These interactions can be time-consuming and emotionally stressful for all parties involved, and some states require agencies to engage with every family that has expressed interest. For example, Florida Administrative Code Rule 65C-16.003 states:
@@
We have also heard families share their frustration with the emotional and time costs that they incurred while participating in a family-driven search system. In a video shared on social media, one adoptive father described his journey that began as a prospective parent engaged with an adoption agency that announced available children and had families respond to express interest in particular children: [^1] https://www.linkedin.com/posts/adoption-share_by-inverting-the-outreach-for-adoption-matching-activity-7131265613947138048-YCM1 (Retrieved 06/17/2025)
@@
In reaction to these challenges, some child welfare agencies have recently sought to improve outcomes by switching to caseworker-driven search (Riley 2019). In this approach, caseworkers sequentially contact specific prospective families to share details about the child. Caseworkers use their expertise to decide which families to contact based on the child’s and the families’ characteristics. Optionally, technological tools may also aid their decisions. This approach removes the burden of engaging a large number of families simultaneously and allows caseworkers to target compatible families for children with very specific needs.
@@
However, even though these are clear advantages of caseworker-driven search, it is important to note that changing the search paradigm may change families’ incentives that influence their interest in different types of children. For a match to occur under either approach, both the caseworker and the prospective family need to agree. This can, for example, mean that if caseworker-driven search leads to a reduction in search costs for families, they may stop being interested in certain children, leaving those children unmatched. As such, it is unclear whether caseworker-driven search leads to more desirable outcomes for the entire population of children requiring adoptive placements.
@@
To assess how different search disciplines affect outcomes, we pursue two complementary approaches in this paper. First, we analyze the strategic behavior of agents under both search disciplines in a game-theoretic search-and-matching model. Second, we provide empirical evidence for the performance of caseworker-driven search compared to classic search approaches by studying outcomes of a real-world child welfare agency that switched its approach in 2018.
@@
Our analytical contributions begin with the introduction of a new search-and-matching model (Section 4), which captures critical features. First, as both approaches to adoption matching are inherently dynamic, we assume that children and families (hereafter referred to as agents) may enter and depart the system at any time. However, to keep the model tractable, we assume the distribution of agent types in the system remains stable over time. Second, we allow for uncertainty regarding whether a child-family pair is compatible. Third, our model captures the heterogeneous preferences of agents, which is a key distinction between our paper and most earlier literature. While we focus specifically on analyzing adoption systems, our work is also relevant to more general search-and-matching theory.
@@
We perform a game-theoretic comparison of the two approaches within our model. In Section 5.2, we establish that (pure strategy) equilibria are guaranteed to exist under both search technologies (i.e., family-driven search and caseworker-driven search). Furthermore, we find that equilibria form a lattice reminiscent of the structure of the set of stable matchings in standard two-sided matching markets. We then present our main theoretical result: Family-driven search equilibrium outcomes can never be Pareto improvements over caseworker-driven search equilibrium outcomes, but there are instances where each caseworker-driven search equilibrium outcome is a Pareto improvement over all family-driven search equilibrium outcomes (Section 6.1). This holds because caseworker-driven search can reduce wasted search efforts. Thus, agents can worry less about accumulating search costs when they express interest in a child. However, because of multiplicity of equilibria and the lattice structure over these equilibria, all children can be strictly better off in either system. The same holds for families. Even when both family-driven search and caseworker-driven search only admit a unique equilibrium each, an agent can be better off in either setting (Section 6.2). This may be surprising, given that caseworker-driven search reduces wasteful search efforts on both sides of the market. We therefore explore the conditions under which caseworker-driven search usually leads to more agents being better off. In Section 7.1, we show that all children will be better off in caseworker-driven search if families are sufficiently impatient. Furthermore, increasing family supply in family-driven search can have a negative effect on children’s utilities, but not in caseworker-driven search (Section 7.2). We find numerically that caseworker-driven search leads to more desirable outcomes for a wide variety of model parameter choices (Section 8).
@@
To supplement these analytical and numerical insights, we empirically analyze case-level data from a technology platform that a multi-county child welfare agency in Florida has been using for the majority of its search efforts since 2018 (Section 9). We compare the outcomes for hundreds of children in need of adoptive placements to a benchmark from statewide case-level data that accounts for a child’s demographic information and disabilities. To measure how child- and case-level factors influence the timing of adoptions in Florida, we fit a Cox proportional hazards model (Cox 1972) to statewide records from the federal Adoption and Foster Care Analysis and Reporting System (AFCARS). We then use this model to predict the adoption outcomes of children listed on the platform if traditional search disciplines were used instead of the agency’s caseworker-driven search approach. This number can be compared to the actual outcome data from the platform. We show that the probability of adoption within three years was 24% higher than the benchmark for the agency’s children. Then, we extend our statistical analysis to directly measure the platform effect using a time-varying Cox model in a conservatively combined dataset. We find a statistically significant 27% improvement in adoption hazard rates under caseworker-driven search compared to statewide case data.
@@ 2 Related Literature
Recently, market designers and operations researchers have shown increased interest in how to best serve historically disadvantaged communities. Researchers have, for example, studied refugee resettlement (Andersson et al. 2018, Bansak et al. 2018, Delacrétaz et al. 2020), the improvement of teacher quality at disadvantaged schools (Combe et al. 2022, 2025), the management of volunteer workforces for non-profit organizations (Berenguer et al. 2023), and allocation of public housing (Arnosti and Shi 2020, Kawasaki et al. 2021, You et al. 2022).
@@
Our paper contributes to the limited literature that studies child welfare systems from operations or market design perspectives — a challenge first articulated by Spindler (1970) and renewed by Slaugh et al. (2025). Slaugh et al. (2016) investigated how the Pennsylvania Statewide Adoption & Permanency Network could utilize a match recommendation tool to improve their process of matching children to prospective parents. Their spreadsheet-based tool can be seen as a simplified version of the previously mentioned data-driven software system. Robinson-Cortés (2019) worked with a foster care data set to analyze placements of children in foster homes. His model predicts that allowing placements across administrative regions would be beneficial for children. MacDonald (2019) studied a dynamic matching problem where children and families can either form reversible matches (foster placements) or irreversible matches (adoptions). In her model, children are heterogeneous in the sense that there are children with disabilities and children without disabilities, while families are homogeneous. To the best of our knowledge, we are the first to formally analyze the economic effects of search and matching in the child welfare domain, while taking into account the full heterogeneity of preferences. Baccara et al. (2014) estimated families’ preferences over children available for adoption from a data set documenting the operations of adoption agencies.
@@
Within the child welfare literature, relatively little research has investigated the effectiveness of search disciplines for children in need of adoptive placements. Some research has reported positive impacts from intensive multi-faceted search efforts by caseworkers: Vandivere et al. (2015) show that children served by skilled recruiters from the Wendy’s Wonderful Kids organization were 1.7 times as likely to have an adoptive placement than a control group in an experiment with over 1,000 children. In a similar context focusing on hard-to-place youth in New York, Feldman et al. (2016) show that a program of enhanced casework improved outcomes for children. The program utilized a variety of channels to promote 88 children and conduct searches. The search methods in both experiments require extensive work from skilled caseworkers funded by grants, while the platform we study provides an example of technology assisting caseworkers. Avery et al. (2009) study national photolisting service AdoptUSKids and use a hazards model to show better outcomes for children based on activity on AdoptUsKids. However, photolistings have drawn increased scrutiny since the early 2000s; Roby and White (2010) describe risks for exploitation and bullying for children publicly listed online.
@@
Even though child adoption matching markets bear similarities to other two-sided matching markets such as centralized labor markets (Roth 1984, 1991) or ride-sharing platforms (Ma et al. 2020), there are important differences that necessitate new models and analyses. Adoption matching is inherently dynamic, and there is no centralized clearinghouse that determines final matches. Matches are only ever proposed, and both sides of the market have to invest search efforts to identify a match candidate.
@@
One matching market similar to adoption from foster care is online dating (Hitsch et al. 2010a, b, Lee and Niederle 2015, Halaburda et al. 2018, Kanoria and Saban 2021). However, most online dating markets are completely decentralized despite their dynamic and recommendation-based features. Individuals in search of a romantic partner can, at any time, decide to browse a dating platform and reach out to other individuals who appeal to them. In contrast, the approaches we analyze in this paper follow a centralized protocol despite the dynamic decentralized search component: Caseworkers perform repeat searches for a family on behalf of children, and they do so in approximately regular time intervals. Caseworkers, therefore, play an essential role throughout the process since they act as an intermediary to protect vulnerable children. Crucially, this introduces an asymmetry, both in the number of agents simultaneously active and the market power of agents on the two sides, that we have not observed in any other previously studied matching market. As a consequence, we develop a new model that allows us to capture the features of the two different search technologies in one model.
Purely random decentralized matching models have been widely studied under search frictions and homogeneous preferences, with transferable utility (Shimer and Smith 2000, 2001, Atakan 2006), and more relevant to our paper, with non-transferable utility (Eeckhout 1999, Chade 2001, Smith 2006). In all of these studies, unlike in our paper, preferences are aligned on each side of the market following a strict order of quality common for each individual on the same side. More recent studies have combined directional search rather than random search as an important feature (Lauermann et al. 2020, Cheremukhin et al. 2020).
Besides the directed versus random search distinction, simultaneous (Stigler 1961) versus sequential (Weitzman 1979) search have also been studied by classic search theory, with Chade and Smith (2006) bridging the gap by characterizing optimal hybrid strategies. Other work extends this analysis to competitive environments, such as labor markets. Albrecht et al. (2006) and Kircher (2009) study workers applying to multiple firms, yielding approximately efficient outcomes through endogenous wage dispersion that coordinates search. More recent work on online platforms also focuses on the trade-off between the two approaches. Honka and Chintagunta (2017) documents that simultaneous consumer search intensifies price competition on insurance aggregators, while Auster et al. (2025) cautions that frictionless simultaneous contacting can exacerbate adverse selection. Chade et al. (2017) and Wright et al. (2021) provide comprehensive surveys of various search models.
The closest to our model are the search theory papers by Adachi (2003) and Lauermann and Nöldeke (2014) on marriage markets, which consider non-symmetric preferences on both sides of the market. However, they only consider a single randomly chosen potential match in each period, while our problem requires that multiple matches with uncertain suitability may be investigated using either caseworker-driven or family-driven search in each period. Similarly, while Immorlica et al. (2023) considers the problem of a centralized platform guiding search through a sequence of match proposals to agents in a two-sided market, they also only consider a single potential match per period. Additionally, they assume that the value for a match is symmetric between both sides. This is unrealistic for an adoption setting, where families’ and children’s desires can be at odds with each other. Consequently, we allow for non-symmetric values.
@@
3 Descriptions of Approaches to Adoption Matching
Based on conversations with caseworkers and managers of a Florida agency that transitioned from a family-driven search to a caseworker-driven search approach, we provide a more detailed introduction to how both approaches function in practice. Although very little has been published about search approaches and their prevalence, subject-matter experts indicate that similar approaches are commonly used nationally, with some variation among states and agencies. In any system, a prospective adoptive family first undergoes an extensive vetting and training process, which usually entails a written home study. Through this report, a caseworker evaluates the family on various dimensions, gathering information from home visits, interviews with family members, third-party sources, and the caseworker’s own judgment. State regulations determine minimum requirements for home studies to assess the suitability of the intended adoptive parents for different types of children. Typically, home studies also include additional information about the family’s environment and preferences. More details on home studies can be found in online Appendix  A.
Regardless of the search approach, agencies may then employ scoring rules to assess family suitability for individual children, sometimes generating and using these scores as part of a recommender system (Slaugh et al. 2016). These tools and others, such as those described by Hanna and McRoy (2011), provide suggestions but are not determinative; all investigation and matching decisions are in the purview of a committee comprising the caseworker, supervisors, and other agency staff.
Once a family has gone through this process, they can apply for adoptive placements for any child for whom the parental rights of the birth family have been terminated. If an interested family is denied a requested placement for non-formal reasons — including the intent to place the child with a different family — Florida Administrative Code 65C-16.005(9) requires that the case be additionally reviewed by a five-person Adoption Applicant Review Committee to ensure a fair evaluation.
The core challenge is making the most suitable families aware of children who need adoptive placements. This is where the agency’s search discipline becomes relevant: identifying and informing suitable families. It should be noted that although there is typically a single agency responsible for finding placements for a child, their search is non-exclusive: any qualified prospective family can, in theory, apply for an adoptive placement and be considered by the agency in the same way as families identified through the agency’s own search. Consequently, a small percentage of children get adopted through other channels, e.g., by friends of their foster (or birth) family.
3.1 Family-driven Search
Traditionally, most agency lies with the families. In a typical family-driven search protocol, the agency emails all approved families when a child becomes eligible for adoption after a judge’s termination of parental rights (TPR) order. The email provides brief details about the child and invites families to express interest. After families have a chance to respond, the caseworker responsible for the child compiles a list of interested families. In most jurisdictions, caseworkers must give serious consideration to every responding family. The caseworker then begins reviewing those families to determine which family is the best suitable match. Obvious mismatches (e.g., a wheelchair-dependent child and a family living in a house not designed to accommodate wheelchairs) can be quickly identified and screened out, but most candidates require a careful review of the home study and follow-up interviews. This process is time-consuming and emotionally taxing for families. If multiple families are suitable, or the decision is complex, the case is referred to a committee, such as the Adoption Applicant Review Committee in Florida, for a final determination. When no family is deemed an adequate fit, the child remains in foster care, and the child’s availability is re-advertised after some time.
3.2 Caseworker-driven Search
In caseworker-driven search, agencies do not issue broad announcements. Instead, once a given child becomes eligible, the child’s caseworker sequentially informs specific families from the approved pool and invites them to apply to adopt the child. Family selection strategies vary across caseworkers. Some rely more on scoring rules and recommender systems, while others may rely on informal networks to prioritize families they know or apply their own heuristics to identify families to consider. If an invited family expresses interest, the investigation mirrors the family-driven search process, including potential committee review. If the caseworker (or the review committee) feels that a family is reasonable but not ideal, they may also invite additional families to apply before coming to a final decision. If the caseworker — who might be managing dozens of cases at different stages of the child welfare system — has exhausted the pool of eligible families without a suitable match emerging, the search is suspended until the family pool has sufficiently refreshed. This means that the child remains in foster care until the search is restarted after some time.